# A Comprehensive Survey on GNN Characterization

### v2

@@ 4.3 Findings on Memory Perspective
Memory Pattern. M-1 The aggregation stage suffers from numerous indirect and irregular memory accesses during both SHoGNN inference [56, 55] and training [62]. This is due to the randomness of neighbor indices for each vertex, leading to extremely high numbers of L2 and L3 cache misses per kilo-instruction. M-2 During SHoGNN inference, memory dependency is the dominant stall in both the scatter and gather substages due to the irregular memory accesses [40, 4]. This aligns with C-4.2.
@@
Data Locality. M-6 SHoGNN training suffers from a low L1 D-cache hit rate across all operations [4]. The average L1 D-cache hit rate is extremely low, at just 15%, which is a primary cause of the stalls noted in C-4.2. M-7 In contrast, the larger L2 cache performs significantly better, with an average hit rate of 70% [4]. M-8 SHoGNN training faces a significant number of divergent load instructions across all operations [4]. An average of 32.5% of divergent load instructions is observed [4]. This percentage is strongly correlated with the low L1 D-cache hit rates in M-4.3.
@@
M-11 During SHoGNN inference, the scatter substage has a high L2 cache hit rate, despite graph processing [57] being known for low spatial data locality [64]. This is because SHoGNNs typically use vertex/edge features of size 32 or more, whereas graph processing like PageRank uses a size of one or two [57]. This finding is consistent with M-4.3 observed during SHoGNN training.
@@
Memory Usage. M-16 SHoGNN training exhibits exponential memory usage complexity with respect to the number of layers [11]. M-17 The peak memory usage during SHoGNN training far exceeds the size of the dataset itself [46]. The high memory usage for intermediate results is the main factor limiting the data scalability of the training and inference [46]. M-18 Most of the memory usage comes from the intermediate results of the aggregation stage [46]. The forward phase generates many intermediate results cached for gradient calculation in the backward phase, increasing memory usage, explaining the high peak memory usage noted in M-4.3.
@@ 4.4 Findings on Communication Perspective
Communication Pattern. CO-1 During mini-batch training of SHoGNN, communication in sampling exhibits a random and fine-grained pattern [26]. On average, 48% of communications are for graph topology, involving indirect pointer-chasing and fine-grained (8-64 byte) transfers [26]. The bandwidth for these small packages is 100$\times$ lower than the peak communication bandwidth, leading to underutilized bandwidth. This observation aligns with M-4.3 regarding memory access patterns.
@@
4.5 Findings on Parallelism Perspective
Parallelism Dependency. P-4 Inter-mini-batch parallelism in SHoGNN training is more complex than in DNN training due to the numerous dependencies between data samples [5]. P-5 Resource contention exists between different training stages when they run simultaneously [31]. Unlike DNN training, GNN training involves complex preprocessing tasks such as sampling and feature collection, which consume CPU, memory, PCIe, and network bandwidth resources. Uncontrolled resource competition can lead to poor performance, with some operations monopolizing resources and blocking others, preventing effective scaling with additional resources [31].
@@ 4.6 Findings on Execution Perspective
Execution Dataflow. E-5 Dataflow exists between stages for each vertex, where the output from the aggregation stage serves as the input for the combination stage [55]. E-6 Performing the combination stage before the aggregation stage can reduce data access and computation in the aggregation stage, as the vertex’s feature vector is typically condensed into a smaller one [55].
@@
E-11 Sampling in mini-batch training is limited by communication during distributed training [26]. As the number of servers increases, sampling throughput scales sub-linearly due to inter-server communication overhead and the random, fine-grained communication pattern revealed by CO-4.4. E-12 Sampling is also a memory usage bottleneck due to the traversal of the entire large-scale graph [26].
@@ 6.4 Findings on Communication Perspective
6.5 Findings on Parallelism Perspective
Parallelism Pattern. P-12 In addition to inter-vertex, intra-vertex, and inter-edge parallelism [55, 53], inter-semantic-graph parallelism exists in the neighbor aggregation stage [58, 15]. This parallelism arises from the independent neighbor aggregation for each semantic graph. As the number of metapaths increases, more opportunities for parallel execution are offered, as indicated by C-6.2 [15].
@@ 7 Characterization Efforts on DHeGNNs
We systematically and comprehensively summarize the findings from the characterization efforts on DHeGNNs.
@@ 8 Summary and Comparison
We systematically summarize and compare the findings across different types of GNNs from six perspectives.
@@ 9.1 Future Directions
Conducting a systematic characterization to compare different types of GNN models is essential for optimizing both hardware designs and software frameworks. This characterization uncovers the commonalities and differences among various GNN models from multiple architectural perspectives. On one hand, it systematically quantifies and compares performance bottlenecks, guiding the design of corresponding software or hardware optimizations. On the other hand, it helps avoid mistakenly transferring potential bottlenecks from one type of model to another.
A promising direction in GNN characterization involves the quantitative analysis of auxiliary aspects, such as security-related issues, including attacks and defenses. Understanding these auxiliary aspects is crucial for ensuring the robustness and security of GNN models in various application scenarios. For example, Wu et al. [49] provided a detailed characterization of defense methods for SHoGNNs on GPUs, highlighting the importance of such analysis for improving model resilience.
@@
A crucial direction is the GNN characterization in industrial scenarios such as recommendation systems, where thousands of GNN tasks are employed for parallel real-time inference. By thoroughly understanding the performance bottlenecks and resource utilization in practical scenarios, we can enhance the efficiency and scalability of GNN models, aiding developers and engineers in formulating more effective optimization strategies. Such characterization not only improves overall system performance but also lays a solid foundation for future advancements in graph-based machine learning technologies.
Establishing comprehensive and standardized benchmarks is crucial for characterizing the performance of all types of GNN models across different hardware platforms and applications. These benchmarks ensure consistency and comparability in research outcomes, help identify performance bottlenecks, and enable targeted optimizations. Similar to MLPerf [36], such benchmarks can drive advancements in GNN research by providing a common ground for performance evaluation.
@@
A promising direction in GNN characterization is the development of fast, automated performance analysis and exploration models. These models streamline the identification and optimization of performance bottlenecks in GNNs through real-time profiling, predictive performance modeling, and automated tuning. By integrating with popular GNN frameworks and ensuring compatibility with diverse hardware platforms, these tools facilitate the resolution of performance bottlenecks in real-time.
Designing hardware accelerators for SHeGNNs and DHeGNNs represents a pivotal future research direction. Current studies have predominantly focused on developing hardware accelerators for SHoGNNs and DHoGNNs, leveraging analyses of performance bottlenecks and optimization opportunities. However, efforts [53, 8, 54, 16] targeting SHeGNNs remain limited, with even fewer studies addressing DHeGNNs. Given the prevalence of static and dynamic heterogeneous graphs in real-world scenarios and their superior capability to represent the relational data in complex systems, there is an urgent need for specialized hardware accelerators.
@@ Acknowledgments
This work was supported by National Key Research and Development Program (Grant No. 2022YFB4501400), the National Natural Science Foundation of China (Grant No. 62202451), CAS Project for Young Scientists in Basic Research (Grant No. YSBR-029), and CAS Project for Youth Innovation Promotion Association.

### v3

@@ 4.4 Findings on Memory Perspective
Memory Pattern. M-1 The aggregation stage suffers from numerous indirect and irregular memory accesses during both SHoGNN inference [54, 53] and training [59]. This is due to the randomness of neighbor indices for each vertex, leading to extremely high numbers of L2 and L3 cache misses per kilo-instruction. M-2 During SHoGNN inference, memory dependency is the dominant stall in both the scatter and gather substages due to the irregular memory accesses [40, 4]. This aligns with C-4.3.
@@
Data Locality. M-6 SHoGNN training suffers from a low L1 D-cache hit rate across all operations [4]. The average L1 D-cache hit rate is extremely low, at just 15%, which is a primary cause of the stalls noted in C-4.3. M-7 In contrast, the larger L2 cache performs significantly better, with an average hit rate of 70% [4]. M-8 SHoGNN training faces a significant number of divergent load instructions across all operations [4]. An average of 32.5% of divergent load instructions is observed [4]. This percentage is strongly correlated with the low L1 D-cache hit rates in M-4.4.
@@
M-11 During SHoGNN inference, the scatter substage has a high L2 cache hit rate, despite graph processing [55] being known for low spatial data locality [61]. This is because SHoGNNs typically use vertex/edge features of size 32 or more, whereas graph processing like PageRank uses a size of one or two [55]. This finding is consistent with M-4.4 observed during SHoGNN training.
@@
Memory Usage. M-16 SHoGNN training exhibits exponential memory usage complexity with respect to the number of layers [11]. M-17 The peak memory usage during SHoGNN training far exceeds the size of the dataset itself [46]. The high memory usage for intermediate results is the main factor limiting the data scalability of the training and inference [46]. M-18 Most of the memory usage comes from the intermediate results of the aggregation stage [46]. The forward phase generates many intermediate results cached for gradient calculation in the backward phase, increasing memory usage, explaining the high peak memory usage noted in M-4.4.
@@ 4.5 Findings on Communication Perspective
Communication Pattern. CO-1 During mini-batch training of SHoGNN, communication in sampling exhibits a random and fine-grained pattern [27]. On average, 48% of communications are for graph topology, involving indirect pointer-chasing and fine-grained (8-64 byte) transfers [27]. The bandwidth for these small packages is 100$\times$ lower than the peak communication bandwidth, leading to underutilized bandwidth. This observation aligns with M-4.4 regarding memory access patterns.
@@ 4.6 Findings on Execution Perspective
Execution Dataflow. E-5 Dataflow exists between stages for each vertex, where the output from the aggregation serves as the input for the combination [53]. E-6 Performing the combination before the aggregation can reduce data access and computation in the aggregation, as the vertex’s feature vector is typically condensed into a smaller one [53].
@@
E-11 Sampling in mini-batch training is limited by communication during distributed training [27]. As the number of servers increases, sampling throughput scales sub-linearly due to inter-server communication overhead and the random, fine-grained communication pattern revealed by CO-4.5. E-12 Sampling is also a memory usage bottleneck due to the traversal of the entire large-scale graph [27].
@@ 6.1 Findings on Time Perspective
6.2 Findings on Parallelism Perspective
Parallelism Pattern. P-12 In addition to inter-vertex, intra-vertex, and inter-edge parallelism [53, 52], inter-semantic-graph parallelism exists in the neighbor aggregation stage [56, 15]. This parallelism arises from the independent neighbor aggregation for each semantic graph. As the number of metapaths increases, more opportunities for parallel execution are offered, as indicated by C-6.3 [15].
@@ 7 Characterization Efforts on DHeGNNs
We systematically and comprehensively summarize the findings from the characterization efforts on DHeGNNs. Due to the limited amount of existing work on DHeGNNs, the summary and discussion of related findings are focused on two key aspects.
@@ 8 Summary and Comparison
We systematically summarize and compare the findings across different GNN types from six key perspectives.
@@
Based on this summary, we present a case study to illustrate how these findings are applied.
A Case Study. We present a case study to demonstrate the practical application of these findings using HyGCN [54], a hardware accelerator tailored for GNNs. First, informed by the finding that the gather and apply_vertex stages dominate execution time ( T-4.1), specialized hardware units are designed to accelerate these critical stages. Second, leveraging the insight that GNNs experience hybrid execution bounds ( E-4.6), an aggregation engine and a combination engine are introduced to address memory-bound and compute-bound challenges, respectively. Finally, based on the observation of inter-stage dataflow between these phases ( E-4.6), a pipeline is established between the two engines. This design leverages dataflow opportunities to reduce unnecessary off-chip memory accesses, thereby enhancing overall efficiency.
Based on this summary, we present several key challenges faced in GNN deployment.
Heterogeneity in Graph Types and Data Sources. GNNs are applied across diverse domains, each with its own unique characteristics. For instance, social networks, biological networks, and knowledge graphs exhibit distinct graph structures and data distributions. The heterogeneity of these data sources makes it challenging to develop a one-size-fits-all deployment strategy. Adapting GNN models to different types of graphs involves customizing the model architecture, data preprocessing pipelines, and training procedures for each domain. Additionally, maintaining consistency, efficiency, and interoperability across these varied graph types and data sources presents a significant challenge in large-scale GNN deployment.
Graph Partitioning and Distribution. Efficient graph partitioning is a critical challenge in deploying GNNs on distributed systems, especially when working with large graphs. The manner in which a graph is divided and distributed across multiple machines has a direct and substantial impact on the performance of the model. Poor partitioning and distribution can lead to increased communication overhead, network congestion, and higher latency due to frequent data exchanges between machines. This not only slows down performance but also reduces scalability.
Scalability with Large-Scale Graphs. A key challenge in deploying GNNs is scaling them to efficiently handle very large graphs. These graphs can consist of billions of vertices and edges, and the compute complexity of GNNs increases significantly as the graph size grows. Scaling GNNs in a distributed environment requires careful attention to memory management, the design of parallelization strategies, and the use of distributed computing frameworks.
Real-Time Processing. Many GNN applications, such as recommendation systems, require real-time or low-latency inference. This necessitates not only fast data retrieval and processing but also the ability to deploy GNNs efficiently in environments where model updates or inferences must occur in real-time. Achieving real-time processing presents a significant challenge, as it demands high throughput and low-latency computation, which becomes particularly difficult when working with large graphs and complex models.
Data Privacy and Security Concerns. When dealing with sensitive data, such as personal information in social networks or medical data in healthcare, data privacy and security become crucial considerations. Ensuring that the distributed processing of graph data does not expose sensitive information is challenging. Techniques such as differential privacy and federated learning are often employed to mitigate privacy concerns, but they introduce additional computational overhead and complexity. Ensuring that sensitive graph data remains secure during processing while maintaining the utility of the model is an ongoing challenge.
@@ 9.1 Future Directions
Characterization on Performance and Scalability. A future direction lies in conducting systematic characterization to compare different types of GNN models, which is vital for optimizing their performance. This effort involves uncovering the shared traits and unique distinctions among various GNN models from diverse architecture perspectives. By systematically identifying and quantifying performance bottlenecks, this direction offers critical insights for designing targeted software and hardware optimizations. Moreover, it helps prevent the inadvertent propagation of inefficiencies from one model type to another. Ultimately, it paves the way for developing more refined and context-aware GNN architectures, ensuring they effectively address challenges specific to their type and deployment scenarios.
A promising direction in GNN characterization is the quantitative analysis of scalability, particularly in the context of large-scale graphs. As GNNs are applied to increasingly massive datasets, understanding how different models scale is essential for optimizing performance. This entails examining how GNNs manage graph size, vertex degree, and edge density while identifying key bottlenecks in data distribution, memory access, and parallelism. By systematically quantifying scalability, this direction provides critical insights for refining GNN architectures, enabling them to handle large graphs more effectively, scale efficiently across distributed systems, and maintain high performance under diverse computational demands.
A promising direction in GNN characterization is the quantitative analysis of graph partitioning techniques, which are pivotal for enhancing the scalability and efficiency of GNN models on large-scale graphs. This analysis examines partitioning methods with respect to graph topology and size, evaluating their impact on both performance and resource utilization. By integrating advanced graph partitioning techniques into the design of GNN architectures, we can achieve improved parallelization, reduced communication costs, and enhanced scalability, thereby optimizing GNN performance in large-scale distributed systems.
Characterization on Cross-domain Integration. A promising direction in GNN characterization is the analysis and optimization of GNN architectures designed for multi-modal graph data, where graph structures integrate diverse formats such as text, images, and time-series data. By assessing the computational and memory requirements specific to multi-modal data, this approach identifies key bottlenecks in feature fusion and cross-modal interaction modeling. These insights facilitate the development of specialized GNN architectures that can efficiently process and integrate heterogeneous data types, thereby broadening their applicability in advanced domains such as multi-modal recommendation systems, autonomous systems, and cross-domain knowledge graphs.
Characterization on Auxiliary Aspects. A key future direction in GNN characterization is the systematic examination of data privacy and security concerns, especially when handling sensitive information such as personal or medical data. This includes characterizing the impact of privacy-preserving techniques, such as differential privacy and federated learning, on GNN performance, particularly regarding computational overhead, scalability, and other related factors. Such studies will provide valuable insights into the trade-offs between ensuring data protection and maintaining model utility.
@@
Characterization on Industrial Scenarios. A crucial direction in GNN characterization is examining GNNs in industrial scenarios, such as recommendation systems, where thousands of tasks are executed for parallel real-time inference. By thoroughly understanding performance bottlenecks and resource utilization in these practical contexts, we can provide developers and engineers with valuable insights for optimization strategies. Additionally, recognizing these real-world constraints helps design GNN architectures that scale effectively while maintaining high performance in production environments. This characterization establishes a strong foundation for future advancements in graph-based machine learning technologies.
@@
Automation and Tooling for Characterization. A promising direction in GNN characterization is the development of fast, automated tools for performance analysis and exploration. These tools enable the identification and optimization of performance bottlenecks in GNNs through real-time profiling, predictive performance modeling, and automated tuning. They accelerate GNN architecture optimization by quickly identifying inefficiencies, enhancing the ability to fine-tune GNN architectures in real-time.
@@ Acknowledgments
This work was supported by National Key R&D Program of China (Grant No. 2023YFB4503500), National Natural Science Foundation of China (Grant No. 62202451 and No. 6230247), CAS Project for Young Scientists in Basic Research (Grant No. YSBR-029), CAS Project for Youth Innovation Promotion Association, Beijing Natural Science Foundation (Grant No. L234078), and Beijing Nova Program (Grant No. 20220484054 and No. 20230484420).